OpenStudy (anonymous):

add 6 and 3/7+4 and 3/4+5 and 4/7

OpenStudy (anonymous):

@welshfella @livias.random

OpenStudy (anonymous):

6 + (3/7 + 4) = 6 + 4 + 3/7 = 10 3/7

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

(5 + 3/4) + 4/7 = 5 + 21/28 + 16/28 = 5 + 37/28 = 6 9/28

OpenStudy (anonymous):

is the answer D

OpenStudy (anonymous):

where are the answer choices?

OpenStudy (anonymous):

I think the earlier person might have done her math wrong.

OpenStudy (anonymous):

4/7 and 3/7 can be combined to equal 1.

OpenStudy (anonymous):

What you are left with is:

OpenStudy (anonymous):

A.15 and 5/14 B.15 and 5/9 C.15 and 3/4 D.16 and 3/4

OpenStudy (anonymous):

6+1+4+5+(3/4)

OpenStudy (anonymous):

This is equal to 16 and 3/4

OpenStudy (anonymous):

so yes it would be D

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

so it is D right

OpenStudy (anonymous):

16 + 3/7 + 9/28 = 16 + 12/28 + 9/28 = 16 + 21/28 = 16 3/4

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@q12157 thanks - I thought they were two separate questions lol

OpenStudy (anonymous):

No worries :3